HELLO,HOW DO YOU CHANGE FROM EXPONENTIAL TO RADICAL FORM AND FROM RADICAL TO EXPONENTIAL. CAN YOU GIVE SOME EXAMPLES? THANKS!!!!
Answer by stanbon(29619)

5^(2/3) = cube root of 5^2
-------------
7^(1/2) = sqrt(7)
-------------------
10^(-5/2) = 1/[sqrt(10^5)]
==================
Cheers,
Stan H.